# High Purcell-enhancement in quantum-dot hybrid circular Bragg grating cavities for GHz-clockrate generation of indistinguishable photons

### v1

@@ Abstract
We present Purcell-enhanced () semiconductor InAs quantum dot radiative lifetimes of ps, enabled by deterministic hybrid circular Bragg gratings (hCBGs). We investigate the benefits of these short times on the two-photon indistinguishability for quasi-resonant and strictly resonant excitation, and observe visibilities at 12.5 ns time delay of consecutively emitted photons. The strongly Purcell-enhanced decay times enable a high degree of indistinguishability for elevated temperatures of up to 30 K, and moreover, allow for excitation of up to 1.28 GHz repetition rate. Our work highlights the prospects of high Purcell-enhanced solid-state quantum emitters for applications in quantum information and technologies operating at GHz clock-rates.
@@ 1 Introduction
Another possibility to minimize the influence of the solid-state environment on the indistinguishability is to decrease the radiative lifetime of the single-photon generating transition, which can be realized by integrating the QD into a photonic cavity and making use of the Purcell effect [21]. Reduced -times increase the two-photon indistinguishability in the presence of inhomogeneous broadening [22], including dephasing from phonons [23, 24], potentially enabling operation without cost-intensive liquid-helium infrastructure.
@@
Short times are furthermore relevant when aiming to excite the QD at high excitation rates. Here, should be ideally considerably shorter than the inverse repetition rate (for example, ), to avoid overlapping of subsequent pulses from following excitation circles, which would require temporal filtering leading to a reduced efficiency due to the discarding of photon events or an increased multi-photon probability. In the past, mostly electrical excitation due to fast possible excitation rates were employed for GHz-driving of In(Ga)As-QD systems, but this limited indistinguishability and a significant amount of multi-photon pulses remained, due to the non-resonant excitation condition and considerable pulse overlap [25, 26, 27, 28]. These limitations have recently been overcome in a work using coherent optical GHz excitation to generate entangled photons from a GaAs-QD [29]. However, no Purcell enhancement was employed here, but rather the faster -time of this material system. Coherent GHz driving of InGaAs QDs with their inherently longer -times has been elusive so far.
@@
To enable Purcell-enhanced -times suited for GHz-operation, close-to-ideal overlapp between the quantum emitter and the cavity mode must be achieved both in the spatial and spectral domain, a challenge hardly mastered to date using non-deterministically fabricated QD devices and resulting in low yield [30]. Photonic cavity systems like micro-pillars [12, 31], photonic crystal cavities [32] or (hybrid) circular Bragg grating cavities (hCBGs) [33, 34, 35] require typically spatial emitter integration precision of well below 50 nm for significant Purcell enhancement.
@@
While several QD-hCBG cavity integration methods are reported as deterministic, this often only implies that a QD is present in the fabricated cavity, but nonetheless significant deviation from theoretically possible Purcell factors is observed, generally attributed to a limited spatial integration precision. Several works have shown experimentally observed Purcell factors of 3-7 [35, 36, 37, 38, 39, 40] with theoretically expected values of 20 from simulations. The highest reported measured Purcell enhancement for QD-hCBG cavities so far are 11.7 [41] and 11.3 [30], respectively, corresponding to -lifetimes of ps in the former case for a GaAs-QD biexciton (XX), and ps for an In(Ga)As XX.
@@ 2 Deterministic fabrication
Before the integration, the QD emission patterns span several m in diameter due to the large distance excited carriers can diffuse in the substrate (Fig. 1(c)). After the deterministic integration, the QD’s emission is confined to the the central disc of the hCBG cavity caused by the interaction with the mode and the finite interaction volume of the insulated disc on the dielectric spacer (Fig. 1(d)). We furthermore observe that the central position of the emission in the CL map after the integration only varies nm, if the spectral region of the QD or of the wetting layer is selected for integration. Thus, rather than the QD-emission profile, the CL-emission map after the integration yields an accurate measure of the hCBG cavity’s center position.
Therefore, comparing the QD’s emission center extracted from CL maps before the fabrication with respect to the specific map’s marker array to the hCBG’s emission center after the fabrication to the same marker array allows for straight forward investigation of the spatial integration precision of the process. The difference between the extracted positions before and after the integration for several hCBG cavities of two exemplary processes and are plotted in Fig. 1(f). Alongside, the spatially-resolved maximum obtained from FEM simulations for for one of the two linear orthogonal polarisations of the hCBG mode is shown. The mean position deviation in and position for Process A (B) are -0.1 nm/13.8 nm (-6.0 nm/30.4 nm) with standard deviations of 10.7 nm/16.0 nm (9.1 nm/14.6 nm). Based on these mean values and standard deviations we conclude that integration precisions 15 nm and accuracies 20 nm can be achieved with our process approach, well within the area of .
@@
The obtained precision- and accuracy-values with the here presented marker-based CL technique are better than the integration performance reported in a recent systematic study that compared marker-based PL- and in-situ CL integration techniques [45]. Reasons for the superior performance of the marker-based CL approach compared to the marker-based PL-approach might be a more accurate determination of the marker positions by the electron beam, but could also originate from the marker-aligned EBL process, e.g. by the considerably higher acceleration voltages used in Ref. [45], which might lead to lower marker contrasts during the alignment, resulting in worse alignment precision. The higher accuracy-values compared to the in situ CL integration technique, which is also carried out at cryogenic temperatures, might originate from the fact, that in case of a present temperature-induced drift during the mapping, the recording of the marker positions can compensate for this drift. It is however also important to note that the sample properties like CL emission pattern and brightness of the QDs can affect the accuracy of determining the QD’s position from the CL maps, and direct comparison between the in-situ CL technique in Ref. [45] and the marker-based CL integration presented here would need to be carried out for the same sample to allow for comparison.
We point out that the mode orientation in the fabricated structures might be rotated in respect to the one shown in Fig. 1(f). The mode orientation here can be regarded as a worst-case-scenario for the observed accuracy spread of the integration. We aim to investigate and reduce the larger integration uncertainty in direction in the future and currently suspect a combination of residual drift and charging effects. We believe the influences of both of these factors to be more pronounced in y-direction, since the CL-mapscans are currently performed as -linescans for subsequent -pixels, meaning that neighbouring -pixels scanned at longer times apart.
@@
With sufficient spatial matching achieved using the marker-based CL-approach described above, we focus now on the spectral overlap between QD and cavity mode. Figure 2(a) shows spectrally resolved data of an exemplary QD-hCBG cavity for varying excitation methods, including the aforementioned CL-spectrum (blue), micro-photoluminescence (PL)-spectra obtained under pulsed laser-excitation with two different excitation energies (green, red), and polarization-resolved reflection spectra (black) obtained by illumination with a broadband white-light-source. The PL and reflection spectra were obtained by transferring the sample to a confocal PL setup and cooling it to 4 K. Detailed information on the optical setup can be found in S.I. S4.
@@
The cavity mode’s spectral position can be extracted from the white-light reflection measurements by fitting the data with a Fano function [46], as indicated in red (blue) for the V- (H-)polarized cavity mode. The extracted V-polarized cavity mode wavelength from the Fano-fit is indicated. Apart from observing the cavity mode dips in the reflected signal, the white light is already sufficient to excite specific states of the embedded QD. The target QD emission line (the positive trion (X+) transition) is indicated. The second line is assigned to the neutral exciton.
@@
The X+ emission line can also be found in the CL spectrum before the hCBG cavity fabrication (blue line in Fig. 2(a)), with additional lines blue-shifted from the trion, which we assign to higher charged states like X2+ [47]. The spectra plotted in green and red show PL-spectra under above-band excitation (excitation wavelength =793 nm) and quasi-resonant excitation (=909.4 nm), respectively. We note, that for this specific wafer material, the neutral states require longer integration times and higher excitation power, to see their signatures in the CL signal, while they appear readily in the off-resonant PL-spectrum. We attribute this to the different excitation mechanisms for CL and PL, with the former exciting (positively) charged states more effectively for the present sample, which stems from the specific given doping background. We refer to Ref. [44], which also showed present positive trion signatures from CL, but with more prominent neutral states. Another indicator for noticeable positive background doping in the given sample is that photoluminescence excitation scans (see. S.I. S5) show that the X+-state is most effectively excited under quasi-resonant conditions.
@@
Fig. 2(d) shows time-resolved measurements of the photon arrival time distribution in semi-logarithmic scaling for spectrally filtered emission from the X+ transitions under pulsed p-shell excitation for QD-hCBGs with varying spectral detuning. The -times are extracted by exponential fits, and are indicated for each curve. The time without a hCBG cavity, i.e. in the surrounding membrane and therefore with a is determined by averaging over several X+ decays measured QDs in the surrounding membrane area under similar excitation conditions to be ps (see. S.I. S5). The extracted -times displayed in Fig. 2 correspond to experimentally measured Purcell enhancements of up to compared to the surrounding membrane in the case of the shortest of ps (red curve), which is still within the resolution of the employed superconducting nanowire detector system, as indicated by the instrument response function (IRF).
The corresponding experimentally determined extracted for different spectral detunings between QD transition and cavity mode of each of the investigated hCBGs is shown in Fig. 1(e). The error margins of the experimentally determined originate from the standard deviation of the distribution of observed planar times for reference. The plotted lines depict FEM simulations for the , with the shaded area in grey indicating the possible variation of because of the range of different -values that were used for the cavity shift, assuming perfect spatial QD-mode matching. Since some of the cavities exhibited reduced Q-factors compared to the simulations (see Fig. 2(b)), the open circles in Fig. 2(e) show the experimentally measured corrected for the reduced Q-factors, resulting in increases between 1% and 45% for better comparison to the simulated Purcell enhancement. The dashed line represents the theoretical Purcell enhancement for a spatial mismatch of nm between QD and hCBG mode. The fact, that the majority of the experimental data lies above this line indicates the high degree of spatial integration precision with the presented approach.
The achieved time in this work is close to the shortest reported In(Ga)As in literature in any kind of photonic resonator to-date, which, to the best of our knowledge,is ps, achieved in a photonic crystal cavity with of 43 [32]. We note that the target of around 30 of the hCBGs presented here could readily increased to 50 and more by modifications of the dielectric spacer and grating parameters [34].
@@
Before turning our focus to the influence of strong Purcell enhancement on the indistinguishability of the emitted photons by the QD-hCBG cavities with varying -times, we investigate the influence of the excitation method on the quantum-optical properties. For details regarding the experimental setup to acquire the following data, including equipment and spectral filtering conditions, we refer once again to S.I. S4.
Figure 3(a) shows the emission spectrum of a QD-hCBG cavity with its X+ emission line at =920.5 nm under quasi-resonant excitation at =891.5 nm at a higher QD-shell (see S.I. S5.) The corresponding time-trace and extracted =54(1) ps ( for this case are shown as an inset. Besides the fast -decay, the time-trace exhibits a second, slower decay of several hundred ps under these excitation conditions.
@@
The identical QD-hCBG cavity can also be excited using strictly resonant excitation (s-shell) of the X+ state to achieve higher TPI-visibilities due to higher degrees of coherence of consecutively emitted photons. The resulting spectrum for a pulse area of is shown in Fig. 3(d). To filter the resonant excitation laser light, a cross-polarization configuration [49] is used.
Note that suppressing the reflected excitation laser to resonantly excite QDs in complex structures such as hCBG cavities is not straightforward due to the increased laser scattering by the photonic structure [50]. Previously, the laser suppression was enhanced by fabricating elliptical hCBG structures [51], while the devices fabricated in this work are symmetrical. Here, the necessary laser suppression is achieved by a low required excitation power due to the high hCBG in-coupling efficiency, so that typical -powers are of the order of 1 nW. The coherent excitation is confirmed by the observation of Rabi rotations (see S.I. S6).
The resulting time-trace for the X+ is shown in the inset. The is slightly reduced to about 42 ps, and also no additional bi-exponential decay is observed in contrast to the quasi-resonant excitation conditions. This indicates that the bi-exponential decay-behaviour stems from a meta-stable decay level which is involved during the relaxation of the charge-carriers into the s-shell before recombination, since it is not observed, if the s-shell is directly populated under strictly resonant excitation.
In contrast to the aforementioned quasi-resonant excitation, the laser pulse was spectrally narrowed-down using a folded 4f pulse-shaper, which caused elongation in the time domain from about 2 ps to about 8 ps. With the temporally longer pulse compared to quasi-resonant excitation, re-excitation starts to limit the observed , which is 0.086(6) under RF excitation, as is shown in Fig. 3(e). Another factor that limits the achievable is insufficient laser suppression, because of scattered resonant laser light scattered by the hCBG cavity, limiting the effectiveness of the spatial suppression pattern.
@@
4 influence on
Extracted TPI visibilities for varying times are displayed in Figure 4(a), for p-shell and s-shell excitation of the specific QD-hCBG system. Note that all the measurements were taken with the identical degree of spectral filtering via a monochromator and following single mode fiber in-coupling, as described above. Both the raw and corrected extracted visibilities (latter with for p-shell and for s-shell, see above) are plotted. The error bars correspond to the measured -values for the specific QD-hCBGs. The dashed lines are -dependent theoretical visibilities based on equation (1) [22].
with the error-function and the factor , including the inhomogeneously broadened QD emission line following a Gaussian distribution with linewidth .
@@
This increased indistinguishability in the presence of decoherence for shortened times is also beneficial, if this decoherence is caused by further inhomogeneous broadening due to phonon interaction [23, 24, 54]. Fig. 4(d) shows extracted TPI visibilities under p-shell excitation for varying degrees of Purcell-enhancement in the temperature range of K. The dotted lines are theoretical TPI visibilities according to
where corresponds to equation (1) and accounts for given dephasing due to the quasi-resonant excitation at given and , and the second factor assumes Markovian-dynamics for the indistinguishabilty [14, 23, 24]. Here, represents the Fourier-limited linewidth and being the broadening of the ZPL due to dephasing by scattering from thermal acoustic photons [24]. According to Ref. [24] we approximate with being the Bose-Einstein function. For the curves in Fig. 4, is set to 3 eV and to 1 meV, similar to previous works [24]. The corresponding -values for each curve at given are between 2.5 and 6 GHz, as obtained by comparison of experimental data and theory in Fig. 4(a).
@@
Fig. 4(d) shows clearly the beneficial effect of strong Purcell enhancement on the photon indistinguishability for elevated temperatures, with still at K for . This temperature is interesting from a technological point-of-view, since it is reachable with compact Stirling cryocoolers, operable without the need of liquid helium cooling [55, 56, 57].
@@
We furthermore point out that changes in due to the temperature induced tuning of the -emission wavelength as well as the hCBG cavity mode also affect the obtained indistinguishabilities. For example, the QD with showed ps at K, and ps at K, reducing from 0.76 to 0.64. The reduction of temperature influence on the indistinguishability at these high is therefore expected to be even higher than what can be seen in Fig. 4, and could be accessed if the QD-hCBG-mode matching could be restored independently from temperature, as was done in previous works e.g. via Stark-tuning [24]. On the other hand, it is worth noting that the spectrally broad Purcell enhancement of hCBG cavities can act as a stabilizing element for temperature shifts, as indicated here by maintaining comparably high indistinguishability values due to low changes in also in the absence of independent tuning possibilities.
5 GHz operation

### v2

@@ Abstract
We present Purcell-enhanced () semiconductor InAs quantum dot decay times of ps, enabled by deterministic hybrid circular Bragg gratings (hCBGs). We investigate the benefits of these short times on the two-photon indistinguishability for quasi-resonant and strictly resonant excitation, and observe visibilities at 12.5 ns time delay of consecutively emitted photons. The strongly Purcell-enhanced decay times enable a high degree of indistinguishability for elevated temperatures of up to 30 K, and moreover, allow for excitation of up to 1.28 GHz repetition rate. Our work highlights the prospects of high Purcell-enhanced solid-state quantum emitters for applications in quantum information and technologies operating at GHz clock-rates.
@@ 1 Introduction
Another possibility to minimize the influence of the solid-state environment on the indistinguishability is to decrease the decay time of the single-photon generating transition, which can be realized by integrating the QD into a photonic cavity and making use of the Purcell effect [40]. Reduced -times increase the two-photon indistinguishability in the presence of inhomogeneous broadening [18], including dephasing from phonons [20, 19], potentially enabling operation without cost-intensive liquid-helium infrastructure.
@@
Short times are furthermore relevant when aiming to excite the QD at high excitation rates. Here, should be ideally considerably shorter than the inverse repetition rate (for example, ), to avoid overlapping of subsequent pulses from following excitation circles, which would require temporal filtering leading to a reduced efficiency due to the discarding of photon events or an increased multi-photon probability. In the past, mostly electrical excitation due to fast possible excitation rates were employed for GHz-driving of In(Ga)As-QD systems, but this limited indistinguishability and a significant amount of multi-photon pulses remained, due to the non-resonant excitation condition and considerable pulse overlap [21, 47, 52]. Optically excited InAs/InP QDs have been reported at GHz clock-rates, but with the typical -times, multiphoton-suppression at GHz was also limited here due to overlap of consecutive pulses [4]. These limitations have recently been overcome in a work using coherent optical GHz excitation to generate entangled photons from a GaAs-QD [24]. However, no Purcell enhancement was employed here, but rather the faster -time of this material system. Coherent GHz driving of InGaAs QDs with their inherently longer -times has been elusive so far.
@@
To enable Purcell-enhanced -times suited for GHz-operation, close-to-ideal overlapp between the quantum emitter and the cavity mode must be achieved both in the spatial and spectral domain, a challenge hardly mastered to date using non-deterministically fabricated QD devices and resulting in low yield [58]. Photonic cavity systems like micro-pillars [53, 15], photonic crystal cavities [29] or (hybrid) circular Bragg grating cavities (hCBGs) [60, 42, 31] benefit from or even require typically spatial emitter integration precision of well below 50 nm for significant Purcell enhancement.
@@
The latter hCBG cavities have gained considerable research interest recently, since they allow for simultaneous out-coupling- and Purcell-enhancement of multiple transitions of the same QD emitter due to their low-Q and broadband cavity properties [58, 31]. They furthermore allow for post-growth optimization of the cavity properties by design modifications [42], which is only possible to a limited extent with e.g micropillar cavities. While several QD-hCBG cavity integration methods are reported as deterministic, this often only implies that a QD is present in the fabricated cavity, but nonetheless significant deviation from theoretically possible Purcell factors is observed, generally attributed to a limited spatial integration precision. Several works have shown experimentally observed Purcell factors of 3-7 [31, 27, 37, 23, 6, 26] with theoretically expected values of 20 from simulations. The highest reported measured Purcell enhancement for QD-hCBG cavities so far are 11.7 [45] and 11.3 [58], respectively, corresponding to -lifetimes of ps in the former case for a GaAs-QD biexciton (XX), and ps for an In(Ga)As XX.
@@ 2 Deterministic fabrication
Before the integration, the QD emission patterns span several m in diameter due to the large distance excited carriers can diffuse in the substrate (Fig. 1(c)). We estimate the error of determining the QD’s position nm from its CL-emission by repeated recording of the same CL-maps, and comparing the extracted positions (see S.I. section S4). After the deterministic integration, the QD’s emission is confined to the the central disc of the hCBG cavity caused by the interaction with the mode and the finite interaction volume of the insulated disc on the dielectric spacer (Fig. 1(d)). We furthermore observe that the central position of the emission in the CL map after the integration only varies nm, if the spectral region of the QD or of the wetting layer is selected for integration. Thus, rather than the QD-emission profile, the CL-emission map after the integration yields an accurate measure of the hCBG cavity’s center position. We can thus estimate the positioning accuracy of the hCBG cavity in the EBL step by comparing the extracted hCBG position to the target value. Two reference processes (see S.I. section S4) yielded an average nm, allowing to estimate the overall integration accuracy as nm.
@@
The obtained accuracy-values with the presented marker-based CL technique here are better than the integration performance reported in a recent systematic study that compared marker-based PL- and in-situ CL integration techniques [33]. Reasons for the superior performance of the marker-based CL approach compared to the marker-based PL-approach might be a more accurate determination of the marker positions by the electron beam, but could also originate from the marker-aligned EBL process, e.g. by the considerably higher acceleration voltages used in Ref. [33], which might lead to lower marker contrasts during the alignment, resulting in worse alignment precision. The higher accuracy-values compared to the in situ CL integration technique, which is also carried out at cryogenic temperatures, might originate from the fact, that in case of a present temperature-induced drift during the mapping, the recording of the marker positions can compensate for this drift. It is however also important to note that the sample properties like CL emission pattern and brightness of the QDs can affect the accuracy of determining the QD’s position from the CL maps, and direct comparison between the in-situ CL technique in Ref. [33] and the marker-based CL integration presented here would need to be carried out for the same sample to allow for comparison.
The estimated integration accuracy using the marker-based CL method in this work is plotted in Fig. 1(f) together with the spatially-resolved maximum obtained from FEM simulations for for one of the two linear orthogonal polarisations of the hCBG mode. We point out that the mode orientation in the fabricated structures might be rotated in respect to the one shown in Fig. 1(f). The mode orientation here can be regarded as a worst-case-scenario, which still predicts that experimental are well within reach for the presented integration approach.
@@
With sufficient spatial matching achieved using the marker-based CL-approach described above, we focus now on the spectral overlap between QD and cavity mode. Figure 2(a) shows spectrally resolved data of an exemplary QD-hCBG cavity for varying excitation methods, including the aforementioned CL-spectrum (blue), micro-photoluminescence (PL)-spectra obtained under pulsed laser-excitation with two different excitation energies (green, red), and polarization-resolved reflection spectra (black) obtained by illumination with a broadband white-light-source. The PL and reflection spectra were obtained by transferring the sample to a confocal PL setup and cooling it to 4 K. Detailed information on the optical setup can be found in S.I. S5.
@@
The cavity mode’s spectral position can be extracted from the white-light reflection measurements by fitting the data with a Fano function [16], as indicated in red (blue) for the V- (H-)polarized cavity mode. The extracted V-polarized cavity mode wavelength from the Fano-fit is indicated. Apart from observing the cavity mode dips in the reflected signal, the white light is already sufficient to excite specific states of the embedded QD. The target QD emission line (the positive trion (X+) transition) is indicated. The second line is assigned to the neutral exciton. The assignment of the QD states is based on excitation power and polarization dependent measurements, as shown in S.I. S6.
@@
The X+ emission line can also be found in the CL spectrum before the hCBG cavity fabrication (blue line in Fig. 2(a)), with additional lines blue-shifted from the trion, which we assign to higher charged states like X2+ [51]. The spectra plotted in green and red show PL-spectra under above-band excitation (excitation wavelength =793 nm) and quasi-resonant excitation (=909.4 nm), respectively. We note, that for this specific wafer material, the neutral states require longer integration times and higher excitation power, to see their signatures in the CL signal, while they appear readily in the off-resonant PL-spectrum. We attribute this to the different excitation mechanisms for CL and PL, with the former exciting (positively) charged states more effectively for the present sample, which stems from the specific given doping background. We refer to Ref. [43], which also showed present positive trion signatures from CL, but with more prominent neutral states. Another indicator for noticeable positive background doping in the given sample is that photoluminescence excitation scans (see. S.I. S7) show that the X+-state is most effectively excited under quasi-resonant conditions.
@@
Fig. 2(d) shows time-resolved measurements of the photon arrival time distribution in semi-logarithmic scaling for spectrally filtered emission from the X+ transitions under pulsed p-shell excitation for QD-hCBGs with varying spectral detuning. The -times are extracted by exponential fits, and are indicated for each curve. The time without a hCBG cavity, i.e. in the surrounding membrane and therefore with a is determined by averaging over several X+ decays measured QDs in the surrounding membrane area under similar excitation conditions to be ps (see. S.I. S7). The extracted -times displayed in Fig. 2 correspond to experimentally measured Purcell enhancements of up to compared to the surrounding membrane in the case of the shortest of ps (red curve), which is still within the resolution of the employed superconducting nanowire detector system, as indicated by the instrument response function (IRF).
The corresponding experimentally determined extracted for different spectral detunings between QD transition and cavity mode of each of the investigated hCBGs is shown in Fig. 2(e). The error margins of the experimentally determined originate from the standard deviation of the distribution of observed planar times for reference. The gray shaded area indicates the different simulated due to the employed varying inner disc size of the hCBG cavity to achieve spectral matching to the QDs, assuming no lateral position deviation. The red solid line corresponds to nm, while the red dashed line is the resulting for nm. All other -dependencies of the fabricated cavities lie between these two curves in the shaded area. Since some of the cavities exhibited reduced Q-factors compared to the simulations (see Fig. 2(b)), the open circles in Fig. 2(e) show the experimentally measured corrected for the reduced Q-factors by , resulting in an increase between 1% and 45% for better comparison to the simulated Purcell enhancement. With the effects of the reduced experimental taken into account, the remaining deviations of the measured Purcell enhancement compared to the simulations are likely to originate from a lateral displacement of the QD from the center of the cavity. The dashed black line represents the theoretical Purcell enhancement for a spatial mismatch of nm between QD and hCBG mode. The fact, that the majority of the experimental data lies above this line indicates the high degree of spatial integration precision with the presented approach.
The achieved -time in this work is close to the shortest reported In(Ga)As -times in literature in any kind of photonic resonator to-date, which, to the best of our knowledge, are ps, achieved in a photonic crystal cavity with of 43 [29], and ps, recently surpassing this -value in a metal-clad nanopillar with of 36 [11]. We note that the target of around 30 of the hCBGs presented here could readily increased to 50 and more by modifications of the dielectric spacer and grating parameters [42].
@@
Before turning our focus to the influence of strong Purcell enhancement on the indistinguishability of the emitted photons by the QD-hCBG cavities with varying -times, we investigate the influence of the excitation method on the quantum-optical properties. For details regarding the experimental setup to acquire the following data, including equipment and spectral filtering conditions, we refer once again to S.I. S5.
Figure 3(a) shows the emission spectrum of a QD-hCBG cavity with its X+ emission line at =920.5 nm under quasi-resonant excitation at =891.5 nm at a higher QD-shell (see S.I. S7.) The corresponding time-trace and extracted =54(1) ps ( for this case are shown as an inset. Besides the fast -decay, the time-trace exhibits a second, slower decay of several hundred ps under these excitation conditions.
@@
The identical QD-hCBG cavity can also be excited using strictly resonant excitation (s-shell) of the X+ state to achieve higher TPI-visibilities due to higher degrees of coherence of consecutively emitted photons. The resulting spectrum for a pulse area of is shown in Fig. 3(d). To filter the resonant excitation laser light, a cross-polarization configuration [38] is used. In addition, we employ weak above band support using a strongly attenuated continuous-wave laser at 730 nm, which was necessary to achieve s-shell excitation for several (but not all) QDs on this sample. We attribute this observation to the stabilization of the charge environment in case of specific quantum emitters.
Note that suppressing the reflected excitation laser to resonantly excite QDs in complex structures such as hCBG cavities is not straightforward due to the increased laser scattering by the photonic structure [56]. Previously, the laser suppression was enhanced by fabricating elliptical hCBG structures [57], while the devices fabricated in this work are symmetrical. Here, the necessary laser suppression is achieved by a low required excitation power due to the high hCBG in-coupling efficiency, so that typical -powers are of the order of 1 nW. The coherent excitation is confirmed by the observation of Rabi rotations as shown in S.I. S8.
The resulting time-trace for the X+ is shown in the inset. The is slightly reduced to about 42 ps, and also no additional bi-exponential decay is observed in contrast to the quasi-resonant excitation conditions. This indicates that the bi-exponential decay-behavior stems from a meta-stable decay level which is involved during the relaxation of the charge-carriers into the s-shell before recombination, since it is not observed, if the s-shell is directly populated under strictly resonant excitation.
In contrast to the aforementioned quasi-resonant excitation, the laser pulse was spectrally filtered using a folded 4f pulse-shaper to improve the spectral matching between laser and X+ emission, resulting in an increase of the laser-pulse width in the time domain from about 2 ps to 8 ps. While limiting to 0.086(6) under RF excitation (cf., Fig. 3, these excitation conditions were evaluated as a good trade-off between re-excitation of the quantum dot and laser-leakage in the collection path, due to scattered laser-light from the fabricated hCBG cavity, limiting the spatial suppression.
@@
3.1 influence on
Extracted TPI visibilities for varying times are displayed in Figure 4(a), for p-shell and s-shell excitation of the specific QD-hCBG system. Note that all the measurements were taken with the identical degree of spectral filtering via a monochromator and following single mode fiber in-coupling, as described above. Both the raw and corrected extracted visibilities (latter with for p-shell and for s-shell, see above) are plotted. The error bars correspond to the measured -values for the specific QD-hCBGs. The dashed lines are -dependent theoretical visibilities based on equation (1) [[18, 36]]:
with the complementary error-function and the factor , including the inhomogeneously broadened QD emission line following a Gaussian distribution with FWHM-linewidth .
@@
This increased indistinguishability in the presence of decoherence for shortened times is also beneficial, if this decoherence is caused by further inhomogeneous broadening due to phonon interaction [20, 19, 8]. Fig. 4(d) shows extracted TPI visibilities under p-shell excitation for varying degrees of Purcell-enhancement in the temperature range of K. Fig. 4(d) shows clearly the beneficial effect of strong Purcell enhancement on the photon indistinguishability for elevated temperatures, with still at K for . This temperature is interesting from a technological point-of-view, since it is reachable with compact Stirling cryocoolers, operable without the need of liquid helium cooling [48, 35, 17].
The solid lines are theoretical TPI visibilities according to
where corresponds to equation (1) and accounts for given dephasing due to the quasi-resonant excitation at given and , and the second factor assumes Markovian-dynamics for the indistinguishability [10, 20, 19]. Here, represents the Fourier-limited linewidth and being the broadening of the ZPL due to dephasing by scattering from thermal acoustic photons [19]. According to Ref. [19] we approximate with being the Bose-Einstein function. For the curves in Fig. 4, is set to 3 eV and to 1 meV, similar to previous works [19]. The corresponding -values for each curve at given are between 2.5 and 6 GHz, as obtained by comparison of experimental data and theory in Fig. 4(a).
@@
Note that the increase in temperature causes an energy shift of X+ and a corresponding change in spectral overlap with the cavity mode. The -time for all four cavities therefore varies for the different temperatures, resulting in a -dependent and thus also in eq. (2). For example, the QD with showed ps at K, and ps () at K, reducing from 0.76 to 0.64. The Purcell enhancement of the two lower cavities was on the other hand increased at K, due to tuning the X+ more into resonance with the respective mode. As shown in S.I. section S9, can be approximated with a linear fit for the considered finite temperature change. The steeper drop of for the two high -cavities is primarily caused by the -dependent . The reduction of temperature influence on the indistinguishability at these high is therefore expected to be even higher than what can be seen in Fig. 4(d), and could be assessed if the QD-hCBG-mode matching could be restored independently from temperature, as was done in previous works e.g. via Stark-tuning [19]. In S.I. section S9, additional calculations are shown if could be kept constant, predicting for the high -device in p-shell excitation. On the other hand, it is worth noting that the spectrally broad Purcell enhancement of hCBG cavities can act as a stabilizing element for temperature shifts, as indicated here by maintaining comparably high indistinguishability values due to low changes in also in the absence of independent tuning possibilities. Last but not least, S.I. section S9 also shows theoretical predictions of the indistinguishability for higher temperature under coherent s-shell excitation. Assuming inhomogenous broadening of MHz as shown in Fig. 4(a), are expected for K.
The observed indistinguishabilities in accordance with theory show the clear benefit of high Purcell enhancement for TPI at elevated temperatures. One should note, that compared to high-Q cavities, e.g., micropillars, the spectrally broad hCBG cavity mode cannot provide a selective enhancement of the ZPL relative to the PSB. This limits the indistinguishability that can be achieved at elevated temperatures without spectral filtering, and the overall efficiency if spectral filtering is applied [20, 25].
3.2 GHz operation